What mass of aluminum is needed to react with 10.0 kg of chromium (III) oxide to produce chromium metal ?


1
Expert's answer
2020-12-17T07:33:58-0500

The balanced chemical equation for this reaction is shown below:

2Al(s) + Cr2O3(s) = 2Cr(s) + Al2O3(s)

n(Cr2O3) = 10000g/152g/mol = 65.8 mol

According to the equation

n(Al) = n(Cr2O3)*2 = 65.8*2 = 131.6 mol

m(Al) = n*M = 131.6mol*27g/mol = 3553 g = 3.55 kg - mass of aluminum, which is needed to react